Complete.
The sum of negative twenty-nine and thirty is negative thirty-three more than a number. What is that number?
717 exceeds twelve times a number by 57. What is the number?
: Complete.
The sum of negative twenty-nine and thirty is negative thirty-three more than a number. What is that number?
717 exceeds twelve times a number by 57. What is the number?

Answer by edjones(2401)

Let x=the number
-29+30=-33+x
-33+x=1
-33+33+x=1+33
x=34
.
Let x=the number
717-12x=57
717-717-12x=57-717
-12x=-660
-12x/-12=-660/-12
x=55
.
Ed